# Deflection angle in the strong deflection limit for static axisymmetric spacetimes: local curvature, matter fields, and quasinormal modes

### v2

@@ V Relation between strong deflection limit coefficients and matter field quantities
Let $T_{(\mu)(\nu)}^{\mathrm{m}}$ denote the value of $T_{(\mu)(\nu)}$ evaluated at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane (i.e., at the unstable photon circular orbit). Then, $V^{\prime\prime}_{\mathrm{m}}$ given in Eq. (55) can be written as
@@
as stated in Eq. (56). In other words, the existence of an unstable photon circular orbit requires that the difference between the tidal components, $E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}$, is sufficiently large compared to the net contribution from the local matter fields $T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}$. When the net matter contribution is negligible (e.g., tipically in vacuum of general relativity), the tidal effects alone determine the orbital instability as
@@
Conversely, if the matter contribution is too large relative to the tidal difference, the condition in Eq. (61) is violated, and the photon circular orbit becomes stable rather than unstable.
@@
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{6}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!3E_{(3)(3)}^{\mathrm{m}}+4\pi\big{(}T_{(0)(0)}^{\mathrm{m}}-T_{(3)(3)}^{\mathrm{m}}+2\>\!T_{(1)(1)}^{\mathrm{m}}+2\>\!T_{(2)(2)}^{\mathrm{m}}\big{)}\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (64)
@@
The coordinate-invariant expressions for the SDL coefficients derived above suggest a potential link between theoretical predictions and observable lensing features near compact objects. If $R_{\mathrm{m}}$ is known, the observationally inferred $\bar{a}$ directly reflects information about the local tidal structure and matter distribution. In combination with accurate modeling of the background geometry, even the subleading term $\bar{b}$ may provide additional insights into the curvature and matter profile near the unstable photon circular orbit.
@@
$\bar{a}$ $=\frac{1}{R_{\mathrm{m}}\sqrt{E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}}}.$ (66)
$\bar{b}$ $=\bar{a}\log\left(\frac{2}{\bar{a}^{2}R_{\mathrm{m}}^{2}E_{(3)(3)}^{\mathrm{m}}}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (67)
@@ VI Strong deflection limit coefficients and Newman–Penrose scalars
These scalars encode the free gravitational field in a coordinate-invariant manner. For a static spacetime, we find that $\Psi_{4}+\Psi_{0}$ is real, $\Psi_{4}-\Psi_{0}$ is purely imaginary, and $\Psi_{1}=\Psi_{3}=0$. Moreover, $\mathbb{Z}_{2}$ symmetry about the $\zeta=0$ plane enforces $\Psi_{0}=\Psi_{4}$, which implies that $\Psi_{0}$ is real on that plane. Under these conditions, the electric part of the Weyl tensor in the chosen tetrad basis are given by
@@
These scalars encode the local matter contributions via the Einstein equations. For a static spacetime, we find that $\Phi_{01}=0$, $\Phi_{12}=0$, and $\Phi_{00}=\Phi_{22}$. Moreover, $\mathbb{Z}_{2}$ symmetry about the $\zeta=0$ plane enforces that $\Phi_{02}$ is real on that plane. Under these conditions, the nonzero NP-Ricci scalars on the $\zeta=0$ plane are given by
@@
We denote by $\Psi_{0}^{\mathrm{m}}$, $\Psi_{2}^{\mathrm{m}}$, and $\Psi_{4}^{\mathrm{m}}$ the Weyl scalars evaluated at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane (i.e., at the unstable photon circular orbit). Similarly, the NP-Ricci scalars at the same location are denoted by $\Phi_{00}^{\mathrm{m}}$, $\Phi_{11}^{\mathrm{m}}$, and $\Phi_{22}^{\mathrm{m}}$. Recalling Eqs. (55), we find that $V^{\prime\prime}_{\mathrm{m}}$ is given by
@@
Since the orbit is unstable (i.e., $V^{\prime\prime}_{\mathrm{m}}<0$ as required in Eq. (56)), it follows that
@@
which encapsulates the necessary balance between the free gravitational field—encoded in $\Psi_{0}^{\mathrm{m}}$ (or $\Psi_{4}^{\mathrm{m}}$)—and the local matter contributions—captured by $\Phi_{00}^{\mathrm{m}}$. Thus, the NP formalism greatly simplifies the criterion for the instability of photon circular orbits.
$\Psi_{0}^{\mathrm{m}}<0,$ (92)
@@
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{1}{\bar{a}^{2}R_{\mathrm{m}}^{2}\big{(}\Psi_{2}^{\mathrm{m}}+\Phi_{11}^{\mathrm{m}}-\mathcal{R}^{\mathrm{m}}/24\big{)}}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (94)
@@ VII Evaluation of the formalism in specific geometries
We illustrate the consistency and physical relevance of the general formalism by evaluating the SDL coefficients in several explicit static, axisymmetric spacetimes, showing how local curvature and matter fields determine light propagation near unstable photon orbits.
@@ VII.1 Zipoy–Voorhees spacetimes
When $\delta=1$, this metric recovers to the Schwarzschild metric.
Solving Eq. (11) yields the coordinate radius of the photon circular orbit as
@@
which implies that $\rho_{\mathrm{m}}$ exists only for $\delta>1/2$. The corresponding circumferential radius of the photon circular orbit, $R_{\mathrm{m}}$, and the critical impact parameter, given in Eq. (10), are
@@
When $\delta=1$ (i.e., the Schwarzschild case), $R_{\mathrm{m}}=3m$ and $b_{\mathrm{c}}=3\sqrt{3}m$. The second derivative of the effective potential at the photon circular orbit is given by
@@
which indicates that the photon circular orbit is unstable.
@@
$R_{\mathrm{m}}^{2}\Psi_{2}^{\mathrm{m}}$ $=\frac{1}{8}\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\delta^{2}},$ (107)
@@
In the static, spherically symmetric limit, we obtain $\bar{a}=1$ and $\bar{b}=\log 6+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi$ (see, e.g., Refs. [4, 5, 35]). This limiting case implies
@@ VII.2 Reissner–Nordström spacetimes
Solving Eq. (11) yields the coordinate radius of the photon circular orbit:
@@
indicating that photon circular orbits exist only when $80\leq Q^{2}\leq 9M^{2}/$ 8. The critical impact parameter (10) is
@@
The second derivative of the effective potential at the photon circular orbit is given by
@@
The condition $V^{\prime\prime}_{\mathrm{m}}<0$ further restricts the electric charge to the range $80\leq Q^{2}<9M^{2}/$ 8.
@@
These results coincide with those obtained in Refs. [9, 10, 35], providing a consistency check for both $\bar{a}$ and $\bar{b}$. When $Q=M$, the coefficients reduce to $\bar{a}=\sqrt{2}$ and $\bar{b}=\sqrt{2}\log 4++I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi$.
@@ VII.3 Majumdar–Papapetrou dihole spacetimes
Solving Eq. (11) yields three distinct solutions for the coordinate radius of the photon circular orbit on the plane $\zeta=0$. Here, we select the branch corresponding to the unstable orbit, which is given by [31, 48, 47]
@@
for $90\leq a<a_{\infty}\equiv 2\sqrt{6}/$ 9 [49]. The corresponding circumferential radius and the critical impact parameter are given by
@@
$b_{\mathrm{c}}$ $=\sqrt{\rho_{\mathrm{m}}}\left(1+\frac{2}{\sqrt{\rho_{\mathrm{m}}+a^{2}}}\right)^{2},$ (127)
@@
We find that $\bar{a}$ coincides with the result in Ref. [31], providing a concrete example that supports our main results. As mentioned in Sec. IV, the contribution from $I_{\mathrm{D}}$ to $\bar{b}$ depends on the choice of $z$. When $a=0$, these coefficients reduce to those for $Q=M$ in Sec. VII.2.
@@ VIII Strong deflection limit and quasinormal modes
A well-established connection exists between QNM frequencies and the instability of photon circular orbits in spherically symmetric spacetimes, particularly in the eikonal limit [36, 37]. The real part of the QNM frequency corresponds to the angular frequency of the photon circular orbit $\Omega_{\mathrm{c}}=1/b_{\mathrm{c}}$, while the imaginary part is determined by the Lyapunov exponent $\lambda_{\mathrm{L}}$ characterizing the instability of the orbit
@@
where $n$ is the overtone number, $l$ is the angular momentum of the perturbation [37]. The Lyapunov exponent is given by (see also Refs.  [50, 51, 52] for recent topics)
@@
which indicates that the Lyapunov exponent appearing in the QNM frequency expression is proportional to the inverse of the SDL coefficient $\bar{a}$. Notably, a similar approach has also been employed on the reflection-symmetric plane of axisymmetric spacetimes, where QNM frequencies are computed using equatorial photon circular orbits. These results suggest a deep geometric link between wave dynamics and lensing in strong gravity.
@@
Our results in the present paper extend the formula (140) to static, axisymmetric spacetimes. Using Eqs. (63) or (93), we recast $\lambda_{\mathrm{L}}$ in the form of Eq. (139) as
@@
These expressions are determined by the coordinate-independent, local curvature and matter field quantities evaluated at the unstable photon circular orbit. This indicates that the damping rate of QNMs is directly linked to the local geometric and matter field properties at the radius of the unstable photon circular orbit.
In the following, we test our geometric expression for the Lyapunov exponent in several static, axisymmetric spacetimes. For each case, we evaluate the NP scalars at the photon circular orbit and confirm agreement with the result obtained from the effective potential. Explicit computations are given in Sec. VII.
@@
These consistent results support the validity of the geometric formula (142) for the Lyapunov exponent in static, axisymmetric spacetimes.
@@ IX Summary and discussion
We study the deflection angle of photons in the SDL in static, axisymmetric spacetimes with reflection symmetry across the equatorial plane. Following the standard method of isolating the logarithmic divergence, we introduce a new variable that allows for a coordinate-invariant formulation of the SDL. We show that the two SDL coefficients, which characterize the logarithmic divergence and its offset, are determined by the second derivative of the effective potential evaluated at the unstable photon circular orbit. This second derivative is further related to local curvature and matter field quantities, yielding a coordinate-invariant expression for these coefficients. This local curvature consists of two distinct contributions: the tidal effects associated with the free gravitational field, described by the electric part of the Weyl tensor, and the matter-induced curvature, encoded in the Einstein tensor, representing the influence of matter fields. We also recast the SDL coefficients in terms of NP scalars, offering a geometric perspective aligned with the null structure of the spacetime. This formulation suggests a promising framework for interpreting observational data in terms of fundamental curvature quantities, particularly through coordinate-invariant combinations of NP Weyl and Ricci scalars. We confirm the consistency of our formulation by applying it to several known spacetimes.
@@
In static, spherically symmetric spacetimes, the SDL coefficients can be expressed solely in terms of the areal radius of the photon sphere and the local energy density and pressure [35]. In contrast, in static, axisymmetric spacetimes, the deflection angle in the SDL is governed not only by local matter fields but also by the free gravitational field encoded in tidal distortions, leading to a more intricate geometric structure. Nevertheless, the leading logarithmic divergence of the deflection angle is still determined entirely by local quantities with clear geometric meaning at the unstable photon circular orbit. This ensures that the formulation retains its coordinate-invariant character.
@@
These results provide a geometrically transparent foundation for connecting theoretical predictions with observations. By expressing the SDL coefficients entirely in terms of coordinate-invariant local quantities, the present formalism establishes a clear link between observable lensing features and the underlying geometry and matter distribution near the unstable photon circular orbit. This, in turn, opens up the possibility of inferring local geometric or matter-field properties—such as tidal forces, energy density, and pressure—from precise measurements of strong gravitational lensing near compact objects.
@@
Through the connection between the QNM and SDL, the present analysis expresses the QNM frequency in terms of local curvature and matter field quantities. This formulation opens up the possibility of probing the local matter and geometry near the unstable photon circular orbit through gravitational wave observations. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [55, 56], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [50, 51].

### v4

@@ V Relation between strong deflection limit coefficients and matter field quantities
Let $T_{(\mu)(\nu)}^{\mathrm{m}}$ denote the value of $T_{(\mu)(\nu)}$ evaluated at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane (i.e., at the unstable circular photon orbit). Then, $V^{\prime\prime}_{\mathrm{m}}$ given in Eq. (55) can be written as
@@
as stated in Eq. (56). In other words, the existence of an unstable circular photon orbit requires that the difference between the tidal components, $E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}$, is sufficiently large compared to the net contribution from the local matter fields $T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}$. When the net matter contribution is negligible (e.g., typically in vacuum of general relativity), the tidal effects alone determine the orbital instability as
@@
Conversely, if the matter contribution is too large relative to the tidal difference, the condition in Eq. (61) is violated, and the circular photon orbit becomes stable rather than unstable.
@@
$\bar{b}$ $=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{6}R_{\mathrm{m}}^{2}\left[\>\!3E_{(3)(3)}^{\mathrm{m}}+4\pi\big(T_{(0)(0)}^{\mathrm{m}}-T_{(3)(3)}^{\mathrm{m}}+2\>\!T_{(1)(1)}^{\mathrm{m}}+2\>\!T_{(2)(2)}^{\mathrm{m}}\big)\>\!\right]\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (64)
@@
The coordinate-invariant expressions for the SDL coefficients derived above indicate a direct link between the curvature and matter quantities and observable lensing features near compact objects. If $R_{\mathrm{m}}$ is known, the observationally inferred $\bar{a}$ directly reflects information about the local tidal structure and matter distribution. In combination with accurate modeling of the background geometry, even the subleading term $\bar{b}$ may provide additional insights into the curvature and matter profile near the unstable circular photon orbit.
@@
$\bar{a}$ $=\frac{1}{R_{\mathrm{m}}\sqrt{E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}}},$ (66)
$\bar{b}$ $=-\bar{a}\log\left(\frac{\bar{a}^{2}}{2}R_{\mathrm{m}}^{2}E_{(3)(3)}^{\mathrm{m}}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (67)
@@ VI Strong deflection limit coefficients and Newman–Penrose scalars
These scalars encode the free gravitational field in a coordinate-invariant manner. The relation between the electric part of the Weyl tensor $E_{(i)(j)}$ and the NP Weyl scalars $\Psi_{A}$ ($A=0,1,\ldots,4$) follows from the standard decomposition of the Weyl tensor in the NP formalism (see Ref. [45]). In a static spacetime with $\mathbb{Z}_{2}$ symmetry, the conditions $\Psi_{1}=\Psi_{3}=0$ and $\Psi_{0}=\Psi_{4}$ (real on the $\zeta=0$ plane) hold, and the nonvanishing components of $E_{(i)(j)}$ on the $\zeta=0$ plane are given by
@@
These scalars encode the local matter contributions via the Einstein equations. By using $G_{ab}=R_{ab}-\frac{1}{2}Rg_{ab}$ and the null tetrad conditions, the Ricci scalar $R$ drops out of $\Phi_{AB}$, so that $R_{ab}$ in $\Phi_{AB}$ can be simply replaced with $G_{ab}$. For a static spacetime, we find that $\Phi_{01}=0$, $\Phi_{12}=0$, and $\Phi_{00}=\Phi_{22}$. Moreover, $\mathbb{Z}_{2}$ symmetry about the $\zeta=0$ plane enforces that $\Phi_{02}$ is real on that plane. Under these conditions with Eqs. (68)–(71), the nonvanishing NP-Ricci scalars on the $\zeta=0$ plane are related to the tetrad components of the Einstein tensor $G_{(\mu)(\nu)}$ as
@@
We denote by $\Psi_{0}^{\mathrm{m}}$, $\Psi_{2}^{\mathrm{m}}$, and $\Psi_{4}^{\mathrm{m}}$ the Weyl scalars evaluated at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane (i.e., at the unstable circular photon orbit). Similarly, the NP-Ricci scalars at the same location are denoted by $\Phi_{00}^{\mathrm{m}}$, $\Phi_{11}^{\mathrm{m}}$, and $\Phi_{22}^{\mathrm{m}}$. Recalling Eqs. (55), we find that $V^{\prime\prime}_{\mathrm{m}}$ is given by
@@
Since the orbit is unstable [i.e., $V^{\prime\prime}_{\mathrm{m}}<0$ as required in Eq. (56)], it follows that
@@
which encapsulates the necessary balance between the free gravitational field—encoded in $\Psi_{0}^{\mathrm{m}}$ (or $\Psi_{4}^{\mathrm{m}}$)—and the local matter contributions—captured by $\Phi_{00}^{\mathrm{m}}$. Thus, the NP formalism greatly simplifies the criterion for the instability of circular photon orbits.
@@
$\bar{b}$ $=-\bar{a}\log\left[\>\!\bar{a}^{2}R_{\mathrm{m}}^{2}\left(\Psi_{2}^{\mathrm{m}}+\Phi_{11}^{\mathrm{m}}-\frac{\mathcal{R}^{\mathrm{m}}}{24}\right)\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (93)
@@ VII Evaluation of the formalism in specific geometries
We illustrate the consistency and physical relevance of the general formalism by evaluating the SDL coefficients in several explicit static and axisymmetric spacetimes, showing how local curvature and matter fields determine light propagation near unstable photon orbits.
@@ VII.1 Zipoy–Voorhees spacetimes
When $\delta=1$, this metric reduces to the Schwarzschild metric.
Solving Eq. (11) yields the coordinate radius of the circular photon orbit as
@@
which implies that $\rho_{\mathrm{m}}$ exists only for $\delta>1/2$. The corresponding circumferential radius of the circular photon orbit, $R_{\mathrm{m}}$, and the critical impact parameter, given in Eq. (10), are
@@
When $\delta=1$ (i.e., the Schwarzschild case), $R_{\mathrm{m}}=3m$ and $b_{\mathrm{c}}=3\sqrt{3}m$. The second derivative of the effective potential at the circular photon orbit is given by
@@
which indicates that the circular photon orbit is unstable.
@@
$R_{\mathrm{m}}^{2}\Psi_{2}^{\mathrm{m}}$ $=\frac{1}{8}\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\delta^{2}}.$ (106)
@@
In the static and spherically symmetric limit, we obtain $\bar{a}=1$ and $\bar{b}=\log 6+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi$ (see, e.g., Refs. [5, 6, 35]). This limiting case implies
@@ VII.2 Reissner–Nordström spacetimes
Solving Eq. (11) yields the coordinate radius of a circular photon orbit:
@@
indicating that the circular photon orbit exists only when $0\leq Q^{2}\leq 9M^{2}/8$. The critical impact parameter (10) is
@@
The second derivative of the effective potential at the circular photon orbit is given by
@@
The condition $V^{\prime\prime}_{\mathrm{m}}<0$ further restricts the electric charge to the range $0\leq Q^{2}<9M^{2}/8$. This confirms that $R_{\mathrm{m}}$ corresponds to the unstable circular photon orbit. Note that for $M^{2}<Q^{2}\leq 9M^{2}/8$ (i.e., the over-extremal regimes), another circular photon orbit appears at $R_{\mathrm{m},-}=(3M-\sqrt{9M^{2}-8Q^{2}})/2$, which corresponds to a stable circular photon orbit. In the extremal case $Q^{2}=M^{2}$, this orbit coincides with the horizon radius and therefore does not represent an independent stable circular photon orbit outside the horizon.
@@
These results coincide with those obtained in Refs. [9, 10, 35], providing a consistency check for both $\bar{a}$ and $\bar{b}$. When $Q=M$, the coefficients reduce to $\bar{a}=\sqrt{2}$ and $\bar{b}=\sqrt{2}\log 4+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi$.
@@ VII.3 Majumdar–Papapetrou dihole spacetimes
Solving Eq. (11) yields three distinct solutions for the coordinate radius of the circular photon orbit on the plane $\zeta=0$. Here, we select the branch corresponding to the unstable orbit, which is given by [33, 48, 47]
@@
for $0\leq a<a_{\infty}\equiv 2\sqrt{6}/9$ [49]. The corresponding circumferential radius and the critical impact parameter are given by
@@
$b_{\mathrm{c}}$ $=\sqrt{\rho_{\mathrm{m}}}\left(1+\frac{2}{\sqrt{\rho_{\mathrm{m}}+a^{2}}}\right)^{2}.$ (126)
@@
We find that $\bar{a}$ coincides with the result in Ref. [33], providing another concrete example that supports our main results. As mentioned in Sec. IV, the contribution from $I_{\mathrm{D}}$ to $\bar{b}$ depends on the choice of $z$. When $a=0$, these coefficients reduce to those for $Q=M$ in Sec. VII.2.
@@ VIII Strong deflection limit and quasinormal modes
A well-established connection exists between QNM frequencies and the instability of circular photon orbits in spherically symmetric spacetimes, particularly in the eikonal limit [36, 37]. The real part of the QNM frequency corresponds to the angular frequency of the circular photon orbit $\Omega_{\mathrm{c}}=1/b_{\mathrm{c}}$, while the imaginary part is determined by the Lyapunov exponent $\lambda_{\mathrm{L}}$ characterizing the instability of the orbit
@@
where $n$ is the overtone number and $l$ is the angular momentum of the perturbation [37]. The Lyapunov exponent is given by
@@
which indicates that the Lyapunov exponent appearing in the QNM frequency expression is proportional to the inverse of the SDL coefficient $\bar{a}$. Notably, a similar approach has also been employed on the reflection-symmetric plane of axisymmetric spacetimes, where QNM frequencies are computed using equatorial circular photon orbits. These results suggest a deep geometric link between wave dynamics and lensing in strong gravity.
@@
Our results in the present paper extend the formula (139) to static and axisymmetric spacetimes. Using Eqs. (63) or (92), we recast $\lambda_{\mathrm{L}}$ in the form of Eq. (138) as
@@
These expressions are determined by the coordinate-independent, local curvature, and matter field quantities evaluated at the unstable circular photon orbit. This indicates that the damping rate of QNMs is directly linked to the local geometric and matter field properties at the radius of the unstable circular photon orbit.
In the following, we test our geometric expression for the Lyapunov exponent in several static and axisymmetric spacetimes. For each case, we evaluate the NP scalars at the circular photon orbit and confirm agreement with the result obtained from the effective potential. Explicit computations are given in Sec. VII.
@@
These consistent results support the validity of the geometric formula (141) for the Lyapunov exponent in static and axisymmetric spacetimes.
@@ IX Summary and discussion
We have studied the deflection angle of photons in the SDL in static and axisymmetric spacetimes with reflection symmetry across the equatorial plane. Following the standard method of isolating the logarithmic divergence, we have introduced a new variable that enabled a coordinate-invariant formulation of the SDL. We have shown that the two SDL coefficients, which characterize the logarithmic divergence and its offset, are determined by the second derivative of the effective potential evaluated at the unstable photon circular orbit. We further have demonstrated that this second derivative can be expressed in terms of local curvature and matter field quantities, yielding a coordinate-invariant expression for these coefficients. This local curvature consists of two distinct contributions: the tidal effects associated with the free gravitational field, described by the electric part of the Weyl tensor, and the matter-induced curvature, encoded in the Einstein tensor, representing the influence of matter fields. We have also recast the SDL coefficients in terms of NP scalars, offering a geometric perspective aligned with the null structure of the spacetime. This formulation suggests a promising framework for interpreting observational data in terms of fundamental curvature quantities, particularly through coordinate-invariant combinations of NP Weyl and Ricci scalars. We have confirmed the consistency of our formulation by applying it to several known spacetimes.
@@
In static and spherically symmetric spacetimes, the SDL coefficients can be expressed solely in terms of the areal radius of the photon sphere and the local energy density and pressure [35]. In contrast, in static and axisymmetric spacetimes, the deflection angle in the SDL is governed not only by local matter fields but also by the free gravitational field encoded in tidal distortions, leading to a more intricate geometric structure. Nevertheless, the leading logarithmic divergence of the deflection angle is still determined entirely by local quantities with clear geometric meaning at the unstable circular photon orbit. This ensures that the formulation retains its coordinate-invariant character.
@@
These results provide a geometrically transparent foundation for connecting theoretical predictions with observations. By expressing the SDL coefficients entirely in terms of coordinate-invariant local quantities, the present formalism establishes a clear link between observable lensing features and the underlying geometry and matter distribution near the unstable circular photon orbit. This, in turn, opens up the possibility of inferring local geometric or matter-field properties—such as tidal forces, energy density, and pressure—from precise measurements of strong gravitational lensing near compact objects. Although the deflection angle itself is not directly measurable, it determines the observable quantities such as the image positions, magnifications, and time delays through the lens equation (see, e.g., Refs. [5, 52]). Therefore, the local curvature expressions of the SDL coefficients obtained in this paper can be directly applied to interpret these measurable lensing features in a coordinate-invariant manner.
@@
Through the connection between the QNM and SDL, we have expressed the QNM frequency in terms of local curvature and matter field quantities. This formulation opens up the possibility of probing the local matter and geometry near the unstable circular photon orbit through gravitational wave observations. Furthermore, our findings may shed light on a universal upper bound on chaos in thermal quantum field theory [53, 54], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [55, 56].
At present, it remains challenging to observationally compare the quantities related to the deflection angle in the SDL with the QNM frequencies because the typical targets of these observations are different: observations of black hole shadows are feasible only for supermassive black holes, whereas QNM frequencies are detected from mergers of stellar-mass binary black holes. However, if the gravitational-wave source itself acts as a lens, it would in principle be possible to observe both phenomena within a single system. Although such a self-lensing scenario is beyond the scope of this work, which considers only static and axisymmetric systems, it would provide a natural setting in which the geometrical and wave aspects of strong gravity could be observed simultaneously.